Circular No. 7993

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002gi, 2002gj, 2002gk, 2002gl, 2002gm Further to IAUC 7977, on behalf of the Supernova Cosmology Project,

R. Knop, Vanderbilt University, reports the discovery of five supernovae at the Cerro Tololo 4-m telescope (+ Mosaic camera) during May 6–8, via subtraction of reference images taken Apr. 15–17 (when the 3σ limit on each magnitude was R = 24.0, except limiting mag I = 24.6 for 2002gi). Spectroscopy and redshifts (given to 0.001 for host galaxies, to 0.01 for supernovae) were obtained during May 8–15 with Gemini North (+ GMOS) on Mauna Kea (2002gm) and with the Very Large Telescope (+ FORS1, FORS2) at Paranal (remaining objects). Photometry of 2002gi was obtained with the Hubble Space Telescope (+ ACS); the other supernovae were followed in R and I bands with the WIYN 3.5-m telescope at Kitt Peak and the 3.5-m Telescopio Nazionale Galileo and 4.2-m William Herschel Telescope at La Palma. Each supernova is within 1′′ of its host galaxy’s center, except for 2002gj (1′′.5 distant).

SN  2002 UT  α_{2000}  δ_{2000}  Mag.  z Type
2002gi May 7 13^{h}57^{m}12.^{s}3 + 4^{o}33′17′′  24.1 I  0.912 Ia
2002gj  7 15 36 25.5 + 9 28 18  22.9 R 0.44  I
2002gk  7 15 37 07.5 + 9 36 19  21.3 R 0.21 Ia
2002gl  7 15 43 24.4 + 7 53 58  23.3 R 0.511 Ia
2002gm  8 15 46 57.0 + 8 12 37  22.3 R 0.49 Ia

SAX J1808.4–3658 C. B. Markwardt, University of Maryland and Goddard Space Flight

Center; J. M. Miller, Harvard-Smithsonian Center for Astrophysics; and R. Wijnands, University of St. Andrews, report a new outburst of the millisecond accreting pulsar SAX J1808.4–3658 (cf. IAUC 7369). In RXTE PCA scans of the Galactic center on Oct. 13.279 UT, an x-ray source whose position is consistent with SAX J1808.4–3658 was detected with a flux of 65 mCrab (2–10 keV). On Oct. 9.5, the PCA flux was < 0.5 mCrab. Two minutes of pointed data detected x-ray pulsations at the frequency 400.953 ± 0.008 Hz (barycentered). This frequency is consistent with the known pulsar spin frequency and orbital Doppler modulation. The RXTE ASM instrument detected the source on Oct. 14.30 and 14.37 at fluxes of 105 and 76 mCrab, respectively. In its 1998 outburst, the source also exhibited fast radio and optical transients, so multiwavelength observing is encouraged.

2002 October 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green